OpenStudy (pphalke):

The scores on a chemistry exam were normally distributed with a mean of 65 and a standard deviation of 9. a. what percent of the students scored above 65? b. what percent of the students scored above 56? c. what percent of the students scored below 47? d. what percent of the students scored between 65 and 74? e. what percent of the students scored between 56 and 83? f. If 1,000 students took the exam and all students with grades 74 and 83 earned a B, how many students earned a B? g. how high must a student score in order to be in the top 2.5% of scores?

OpenStudy (welshfella):

you use the same table as the previous post. Z for the first score 65 = 65-65 / 9 = 0

OpenStudy (welshfella):

what is the reading for 0?

OpenStudy (pphalke):

.50000

OpenStudy (welshfella):

right so that simply means that 1 - 0.5 = 0.5 = 50% is the answer

OpenStudy (welshfella):

Z score for b is 56-65 / 9 = -1

OpenStudy (pphalke):

so I look for the -1 reading?

OpenStudy (welshfella):

yea

OpenStudy (pphalke):

my reading shows .15866

OpenStudy (welshfella):

right remember this represents the area to the left of the Z score. But we require the area to the RIGHT becuase we have scores of above 56

OpenStudy (welshfella):

NOw the area of the whole graph is 1 so the answer is 1 - 0.15866 do you understand that ?

OpenStudy (pphalke):

yes I do understand I was thinking of it myself

OpenStudy (welshfella):

good

OpenStudy (welshfella):

c is a direct reading because our table gives values below the Z score . so for scores < 47 the Z score is 47-65 / 9 = -2 just read -2 off the table.

OpenStudy (welshfella):

sorry gtg right now